# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v1

@@ 4. The dependent self-affine case
Let $f\in\mathcal{S}(\mathbb{R}^{2})$. If $f(K)\subset K$, then we can find $g\in\mathcal{S}(R^{2})$ and a closed set $C\subset[0,2]$ with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ such that $g(K)\subset C\times\mathbb{R}$. Moreover, if $f$ is oblique, then $g$ can be picked oblique as well.
@@
Let $\lambda$ be the contraction ratio of $f$. For $t\geq 1$, write
@@
Since the contraction ratio of $f^{t}$ is $\lambda^{t}$, $K\subset[0,1]^{2}$ and $f^{t}(K)\subset K$, $f^{t}(K)$ is contained in at most $9$ open squares of side length $m^{-pk_{t}}=n^{-qk_{t}}$. More precisely, there are $x\in\frac{\mathbb{N}}{n^{qk_{t}}},y\in\frac{\mathbb{N}}{m^{pk_{t}}}$ such that
@@
Denote by $h_{t}$ the homothety sending $(0,1)^{2}$ to the square $(x,x+n^{-qk_{t}})\times(y,y+m^{-pk_{t}})=:Q^{\prime}_{t}$, the “left bottom square” of $Q_{t}$. Note that for all $t\geq 1$,
@@
The contraction ratio of $h_{t}^{-1}f^{t}$ equals $n^{qk_{t}}\lambda^{t}$, which is at most $n^{qk_{t}}\cdot m^{-pk_{t}}/m=m^{-1}$ and at least $n^{qk_{t}}\cdot m^{-p(k_{t}+1)}/m=m^{-p-1}$. So the sequence $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$ has accumulation points in $\mathcal{S}(\mathbb{R}^{2})$ and we may assume that $h_{t}^{-1}f^{t}$ converges to some $g$.
Moreover, write $y=\sum_{k=1}^{pk_{t}}y_{k}m^{-k}$, where $\{y_{k}\}_{k=1}^{\infty}\subset J$. Since $p>q$, the collection of all level-$pk_{t}$ rectangles in $\overline{Q^{\prime}_{t}}$ is
$\mathcal{R}_{t}:=\Big{\{}\Big{[}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}},\Big{(}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}}\Big{)}+n^{-pk_{t}}\Big{]}\times[y,y+m^{-pk_{t}}]:(x_{k},y_{k})\in\Lambda\Big{\}}.$
Writing $\sigma$ to be the map sending $y$ to $my\pmod{1}$, we have
@@
$\subset h_{t}^{-1}\Big{(}\bigcup\mathcal{R}_{t}\Big{)}$
$\subset\bigcup_{\begin{subarray}{c}(x_{k},y_{k})\in\Lambda,\\ qk_{t}+1\leq k\leq pk_{t}\end{subarray}}\Big{[}\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-pk_{t}}},n^{-(p-q)k_{t}}+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-pk_{t}}}\Big{]}\times\mathbb{R}$
(4.2) $=K^{\sigma^{qk_{t}}(y)}_{(p-q)k_{t}}\times\mathbb{R},$
where $K^{\sigma^{qk_{t}}(y)}_{(p-q)k_{t}}$ is defined in (2.5).
Let $z$ be an accumulation point of $\{\sigma^{qk_{t}}(y)\}_{t=1}^{\infty}$ and write $z=\sum_{k=1}^{\infty}z_{k}m^{-k}$, where $z_{k}\in J$. Using Cantor’s diagonal argument, it is not hard to pick $z$ properly so that for all $j\geq 1$, there is some $t$ (depending on $j$) with
(4.3) $z_{1}=y_{qk_{t}+1},z_{2}=y_{qk_{t}+2},\ldots,z_{j}=y_{qk_{t}+j},$
i.e., the first $j$ terms of the $m$-ary expansion of $z$ and $\sigma^{qk_{t}}(y)$ are identical. We claim that
@@
If the left hand side is empty then there is nothing to prove. Suppose it is non-empty and (4.4) does not hold. Then there is some $b\in g(K)\cap(0,1)^{2}$ such that
@@
where $\operatorname{dist}(A,B):=\inf\{|a-b|:a\in A,b\in B\}$ is the usual distance. Let $M$ be a large integer with $n^{-M}<\frac{\delta}{4}$. Pick $t$ so large that $d_{H}(g(K)\cap[0,1]^{2},h_{t}^{-1}f^{t}(K)\cap[0,1]^{2})<\frac{\delta}{4}$ (where $d_{H}$ is the Hausdorff distance, see [8]), $n^{-(p-q)k_{t}}<\frac{\delta}{4}$ and (4.3) holds for $j=M$. The first requirement gives us a $b_{1}\in h_{t}^{-1}f^{t}(K)\cap[0,1]^{2}$ with $|b_{1}-b|<\frac{\delta}{4}$. By (4.3),
$\operatorname{dist}(K^{\sigma^{qk_{t}}(y)}_{(p-q)k_{t}},K^{z})\leq n^{-\min\{(p-q)k_{t},M\}}.$
Hence
@@
$\leq\operatorname{dist}(K^{\sigma^{qk_{t}}(y)}_{(p-q)k_{t}},K^{z})$ (by (4.2))
$\leq n^{-\min\{(p-q)k_{t},M\}}<\delta/4.$
@@
Applying the same argument to the other eight squares in $Q_{t}$, we find nine points $z_{0,0}:=z,z_{0,1},\ldots,z_{2,2}\in\pi_{2}(K)$ such that
(4.5) $g(K)\subset\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+(i,j))\times\mathbb{R},$
where $K^{z_{i,j}}+(i,j)$ is the translation of $K^{z_{i,j}}$ by $(i,j)$. Write $C:=\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+i)$. By (2.6), $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$.
It remains to prove that if $f$ is oblique then $g$ can be picked oblique as well. Write $O$ to be the orthogonal part of $f$. If $\{O^{t}\}_{t=1}^{\infty}$ is a finite collection, then every element in it appears infinitely many times. Picking an oblique transformation $\widetilde{O}\in\{O^{t}\}_{t}$ and a subsequence $t_{k}$ so that $O^{t_{k}}=\widetilde{O}$ for all $k$, we can consider $f^{t_{k}}$ instead of $f^{t}$ in the above arguments instead and get an oblique $g$ because $g$, as the limit of $h_{t}^{-1}f^{t}$, has $\widetilde{O}$ as its orthogonal part. If $\{O^{t}\}_{t=1}^{\infty}$ is an infinite collection, then since an orthogonal $2\times 2$ matrix is either a rotation or a reflection, $O$ must be an irrational rotation (i.e., the rotation angle is an irrational multiple of $\pi$) and hence every rotation matrix is an accumulation point of $\{O^{t}\}_{t=1}^{\infty}$. Then one just picks a subsequence $\{O^{t_{k}}\}_{k}$ so that $O^{t_{k}}$ converges to an oblique transformation, which is the orthogonal part of $g$. This completes the proof. ∎
@@
In the above proof, if we apply Baire’s theorem to (4.5), some $(K^{z_{i,j}}+(i,j))\times\mathbb{R}$ should contain an interior part of $g(K)$. In particular, there is an integer $k$ and a level-$k$ rectangle $R$ such that $g\varphi_{R}(K)\subset(K^{z_{i,j}}+(i,j))\times\mathbb{R}$. As a consequence,
$g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z_{i,j}}+(i,j))\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
We also remark that the integer $k_{t}$ defined in (4.1) obviously equals $\lfloor(\frac{t\log\lambda}{\log m}-1)/p\rfloor$. So if $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, then $\frac{\log\lambda}{\log m}\notin\mathbb{Q}$ and hence $\{n^{qk_{t}}\lambda^{t}\}_{t=1}^{\infty}$ is dense in an open subinterval of $(0,1)$. Since the first part of Lemma 4.1 holds for all accumulation point of $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$, one can find such an accumulation point $\widetilde{g}$ so that $\widetilde{g}(K)\subset C_{\widetilde{g}}\times\mathbb{R}$, where $C_{\widetilde{g}}$ is a closed set with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ (just as $C$ in Lemma 4.1), and the contraction ratio of $\widetilde{g}$ is an irrational power of $n$.
@@
By Remark 3.3 and Proposition 2.2, it suffices to consider when $N\leq n-1$ and #$\#J\leq m-1$. Suppose on the contrary that $f$ is oblique. Pick $g,C$ be as in Lemma 4.1 with $g$ oblique and denote by $\rho$ the contraction ratio of $g$. If $N=1$, then it is not hard to see that every horizontal slice of $K$ is a finite set. From the form of $C$ (recall (4.5)), $C$ is a finite set. But by Proposition 2.1 and the obliqueness of $g$, $\pi_{1}(g(K))$ is an infinite set. This contradicts $\pi_{1}(g(K))\subset\pi_{1}(C\times\mathbb{R})=C$.
Now suppose $N\geq 2$. By Remark 4.2, we can find $z\in\pi_{2}(K)$, $c\in\mathbb{R}^{2}$, $k\geq 1$ and a level-$k$ rectangle $R$ such that
@@
Next, since $h_{z}(K^{z})\subset K^{z}$, we have by Corollary 2.7 two points $a,b\in K^{z}$ such that $a,h(a)$ (resp. $b,h(b)$) are both left (resp. right) isolated points of $K^{z}$. Since $h_{z}(K^{z})\subset K^{z}$ is simply a one-dimensional realization of the inclusion $g\varphi_{R}(K^{z}\times\{z\})\subset L_{z}\cap((K^{z}+c)\times\mathbb{R})$, this in turn gives us two points $\widetilde{a},\widetilde{b}\in g\varphi_{R}(K^{z}\times\{z\})$ and some $\delta>0$ such that both of the tubes $(\pi_{1}(\widetilde{a})-\tfrac{\delta v_{1}}{|v|},\pi_{1}(\widetilde{a}))\times\mathbb{R}$ and $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\frac{\delta v_{1}}{|v|})\times\mathbb{R}$ do not intersect $L_{z}\cap((K^{z}+c)\times\mathbb{R})$. Moreover, writing $R_{t},R^{\prime}_{t}\subset R$ to be the level-$t$ rectangles such that $\widetilde{a}\in g\varphi_{R_{t}}(K)$ and $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, respectively, we have (again by Corollary 2.7) $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all $t\geq k_{0}$ for some $k_{0}$. See Figure 4 for an illustration.
@@
Figure 4. An illustration of $\widetilde{a},\widetilde{b},R_{t}$ and $R^{\prime}_{t}$, where $(K^{z}+c)\times\mathbb{R}$ is supported in the shaded region. (color online)
Pick $x,x^{\prime}\in K$ with $\pi_{2}(x)=\max\pi_{2}(K)$ and $\pi_{2}(x^{\prime})=\min\pi_{2}(K)$. For convenience, let $e_{1}:=(1,0)$ and $e_{2}:=(0,1)$. Since $\widetilde{a}\in g\varphi_{R_{t}}(K)$, $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, $x$ can be written as
@@
So at least one of $d_{t},-d^{\prime}_{t}$ is strictly larger than $\beta/3$. Writing $e_{1,1}:=\pi_{1}(g(e_{1}))$ and $e_{2,1}:=\pi_{1}(g(e_{2}))$, we have by the obliqueness of $g$ that $e_{1,1},e_{2,1}\neq 0$. We distinguish four simple cases.
Case 1: $d_{t}>\beta/2$ and $e_{2,1}>0$
@@
$g\varphi_{R^{\prime}_{t}}(x)=\widetilde{b}+g\varphi_{R_{t}}(\gamma_{t}e_{1})+g\varphi_{R_{t}}(d_{t}e_{2}).$
Since $d_{t}>\beta/2$, $n>m$ and $|\gamma_{t}|\leq 1$, $0<\rho n^{-t}\gamma_{t}e_{1,1}+\rho m^{-t}d_{t}e_{2,1}<\frac{\delta v_{1}}{|v|}$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x))=\pi_{1}(\widetilde{b})+\rho n^{-t}\gamma_{t}e_{1,1}+\rho m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\tfrac{\delta v_{1}}{|v|}).$
In other words, $g\varphi_{R^{\prime}_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\frac{\delta v_{1}}{|v|})\times\mathbb{R}$. But we have seen that this tube does not intersect $L_{z}\cap((K^{z}+c)\times\mathbb{R})\supset g\varphi_{R}(K)\ni g\varphi_{R^{\prime}_{t}}(x)$, which leads to a contradiction.
Case 2: $d_{t}>\beta/2$ and $e_{2,1}<0$
@@
$g\varphi_{R_{t}}(x)=\widetilde{a}+g\varphi_{R_{t}}(c_{t}e_{1})+g\varphi_{R_{t}}(d_{t}e_{2}).$
Since $d_{t}>\beta/2$, $n>m$ and $|c_{t}|\leq 1$, we have $-\frac{\delta v_{1}}{|v|}<\rho n^{-t}c_{t}e_{1,1}+\rho m^{-t}d_{t}e_{2,1}<0$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R_{t}}(x))=\pi_{1}(\widetilde{a})+\rho n^{-t}c_{t}e_{1,1}+\rho m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\tfrac{\delta v_{1}}{|v|},\pi_{1}(\widetilde{a})).$
In other words, $g\varphi_{R_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{a})-\tfrac{\delta v_{1}}{|v|},\pi_{1}(\widetilde{a}))\times\mathbb{R}$. Again, we have seen that this tube does not intersect $L_{z}\cap((K^{z}+c)\times\mathbb{R})\supset g\varphi_{R}(K)\ni g\varphi_{R_{t}}(x)$, which leads to a contradiction.
Case 3: $d^{\prime}_{t}>\beta/2$ and $e_{2,1}>0$
@@
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x^{\prime}))=\pi_{1}(\widetilde{b})+\rho n^{-t}\gamma^{\prime}_{t}e_{1,1}+\rho m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\tfrac{\delta v_{1}}{|v|}),$
which will lead to a contradiction.
Case 4: $d^{\prime}_{t}>\beta/2$ and $e_{2,1}<0$
@@
$\pi_{1}(g\varphi_{R_{t}}(x^{\prime}))=\pi_{1}(\widetilde{a})+\rho n^{-t}c^{\prime}_{t}e_{1,1}+\rho m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\tfrac{\delta v_{1}}{|v|},\pi_{1}(\widetilde{a})),$
which will lead to a contradiction. ∎
Combining with the above non-obliqueness result and Proposition 2.6, we can quickly prove Theorem 1.5.

### v2

@@ 4. The dependent self-affine case
Let $f\in\mathcal{S}(\mathbb{R}^{2})$. If $f(K)\subset K$, then we can find $g\in\mathcal{S}(R^{2})$ and a compact set $C\subset\mathbb{R}$ with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ such that $g(K)\subset C\times\mathbb{R}$. Moreover, if $f$ is oblique, then $g$ can be picked oblique as well.
@@
Let $\lambda$ be the contraction ratio of $f$. For all positive integers $t$ with $\lambda^{t}\leq m^{-1}$, write
@@
Since the contraction ratio of $f^{t}$ is $\lambda^{t}$, $K\subset[0,1]^{2}$ and $f^{t}(K)\subset K$, $f^{t}(K)$ is contained in at most $9$ squares of side length $m^{-pk_{t}}=n^{-qk_{t}}$. More precisely, there are $x\in\frac{\mathbb{N}}{n^{qk_{t}}},y\in\frac{\mathbb{N}}{m^{pk_{t}}}$ such that
@@
Denote by $h_{t}$ the homothety sending $(0,1)^{2}$ to the square $(x,x+n^{-qk_{t}})\times(y,y+m^{-pk_{t}})=:Q^{\prime}_{t}$, the “left bottom square” of $Q_{t}$. Note that for all $t$,
@@
The contraction ratio of $h_{t}^{-1}f^{t}$ equals $n^{qk_{t}}\lambda^{t}$, which is at most $n^{qk_{t}}\cdot m^{-pk_{t}-1}=m^{-1}$ and at least $n^{qk_{t}}\cdot m^{-p(k_{t}+1)-1}=m^{-p-1}$. So the sequence $\{h_{t}^{-1}f^{t}\}_{t}$ has accumulation points in $\mathcal{S}(\mathbb{R}^{2})$. For simplicity, assume that $h_{t}^{-1}f^{t}$ converges to some $g$.
Moreover, write $y=\sum_{k=1}^{pk_{t}}y_{k}m^{-k}$, where $y_{k}\in\{0,\ldots,m-1\}$. Since $p>q$, the collection of all level-$pk_{t}$ rectangles in $\overline{Q^{\prime}_{t}}$, denoted by $\mathcal{R}_{t}$, is (if non-empty)
$\Big{\{}\Big{[}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}},\Big{(}x+\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k}}\Big{)}+n^{-pk_{t}}\Big{]}\times[y,y+m^{-pk_{t}}]:(x_{k},y_{k})\in\Lambda\Big{\}}.$
Then we have
@@
(4.2) $\subset h_{t}^{-1}\Big{(}\bigcup\mathcal{R}_{t}\Big{)}$
$\subset\bigcup_{\begin{subarray}{c}(x_{k},y_{k})\in\Lambda,\\ qk_{t}+1\leq k\leq pk_{t}\end{subarray}}\Big{[}\sum_{k=qk_{t}+1}^{pk_{t}}\frac{x_{k}}{n^{k-qk_{t}}},n^{-(p-q)k_{t}}+\sum_{k=qk_{t}+1}^{k_{t}}\frac{x_{k}}{n^{k-qk_{t}}}\Big{]}\times\mathbb{R}$
(4.3) $=:E_{t}\times\mathbb{R}.$
Suppose $g(K)\cap(0,1)^{2}\neq\varnothing$. Then $h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}\neq\varnothing$ for all large $t$. For each of these $t$, note that $\mathcal{R}_{t}\neq\varnothing$ (recall (4.2)) and write $\bm{j}_{t}:=y_{qk_{t}+1}\cdots y_{pk_{t}}$, which is a word of length $(p-q)k_{t}\to\infty$ as $t\to\infty$. Using Cantor’s diagonal argument, it is not hard to find an infinite word $\bm{j}=j_{1}j_{2}\cdots\in J^{\infty}$ such that for all $M\geq 1$, there is a large $t$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$, where $\bm{j}_{t}\wedge\bm{j}$ denotes their longest common prefix and ||$\cdot$|| ⋅ | stands for the length. Let $z:=\sum_{k=1}^{\infty}j_{k}m^{-k}$.
We claim that
@@
If not, then there is some $b\in g(K)\cap(0,1)^{2}$ such that
@@
where $\operatorname{dist}(A,B):=\inf\{|a-b|:a\in A,b\in B\}$ is the usual distance. Let $M$ be a large integer with $n^{-M}<\frac{\delta}{4}$. Pick $t$ so large that the following conditions hold:
  1. (1)
    $d_{H}(g(K)\cap(0,1)^{2},h_{t}^{-1}f^{t}(K)\cap(0,1)^{2})<\frac{\delta}{4}$ (where $d_{H}$ is the Hausdorff distance, see [8]);
  2. (2)
    $(p-q)k_{t}\geq M$;
  3. (3)
    there is some $\bm{j}_{t}$ such that $|\bm{j}_{t}\wedge\bm{j}|\geq M$.
The first condition gives us a $b_{1}\in h_{t}^{-1}f^{t}(K)\cap(0,1)^{2}$ with $|b_{1}-b|<\frac{\delta}{4}$. Since $|\bm{j}_{t}\wedge\bm{j}|\geq M$, it is not hard to see from the definition of $E_{t}$ and $\bm{j}_{t}$ that $\operatorname{dist}(E_{t},K^{z})\leq n^{-M}$. Hence
@@
$\leq\operatorname{dist}(E_{t},K^{z})$ (by (4.3))
$\leq n^{-M}<\delta/4.$
@@
Applying the same argument to the other eight squares in $Q_{t}$, we find at most nine points $z_{0,0}:=z,z_{0,1},\ldots,z_{2,2}\in\pi_{2}(K)$ such that
(4.5) $g(K)\subset\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+i)\times\mathbb{R}.$
Write $C:=\bigcup_{0\leq i,j\leq 2}(K^{z_{i,j}}+i)$. By (2.6), $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$.
It remains to prove that if $f$ is oblique then $g$ can be picked oblique as well. Write $O$ to be the orthogonal part of $f$. If $\{O^{t}\}_{t=1}^{\infty}$ is a finite collection, then every element in it appears infinitely many times. Picking an oblique transformation $\widetilde{O}\in\{O^{t}\}_{t}$ and a subsequence $t_{k}$ so that $O^{t_{k}}=\widetilde{O}$ for all $k$, we can consider $f^{t_{k}}$ instead of $f^{t}$ at the beginning of the above arguments and get an oblique $g$ because $g$, as the limit of $h_{t}^{-1}f^{t}$, has $\widetilde{O}$ as its orthogonal part. If $\{O^{t}\}_{t=1}^{\infty}$ is an infinite collection, then since a $2\times 2$ orthogonal matrix is either a rotation or a reflection, $O$ must be an irrational rotation (i.e., the rotation angle is an irrational multiple of $\pi$) and hence every rotation matrix is an accumulation point of $\{O^{t}\}_{t=1}^{\infty}$. Then one just picks a subsequence $\{O^{t_{k}}\}_{k}$ so that $O^{t_{k}}$ converges to an oblique transformation, which, similarly as before, should be the orthogonal part of $g$. This completes the proof. ∎
@@
In the above proof, if we apply Baire’s theorem to (4.5), some $(K^{z_{i,j}}+i)\times\mathbb{R}$ should contain an interior part of $g(K)$. In particular, there is an integer $k$ and a level-$k$ rectangle $R$ such that $g\varphi_{R}(K)\subset(K^{z_{i,j}}+i)\times\mathbb{R}$. As a consequence,
(4.6) $g\varphi_{R}(K^{y}\times\{y\})\subset(K^{z_{i,j}}+i)\times\mathbb{R},\quad\forall y\in\pi_{2}(K).$
We also remark that the integer $k_{t}$ defined in (4.1) equals $\lfloor(-\frac{t\log\lambda}{\log m}-1)/p\rfloor$. So if $\frac{\log\lambda}{\log n}\notin\mathbb{Q}$, then $\frac{\log\lambda}{\log m}\notin\mathbb{Q}$ and hence $\{n^{qk_{t}}\lambda^{t}\}_{t}$ is dense in an open subinterval of $(0,1)$. Since the proof of Lemma 4.1 works for all convergent subsequence of $\{h_{t}^{-1}f^{t}\}_{t=1}^{\infty}$, one can find such an accumulation map $\widetilde{g}$ so that $\widetilde{g}(K)\subset C_{\widetilde{g}}\times\mathbb{R}$, where $C_{\widetilde{g}}$ is a closed set with $\operatorname{dim_{H}}C\leq\frac{\log N}{\log n}$ (just as $C$ in Lemma 4.1), and the contraction ratio of $\widetilde{g}$ is an irrational power of $n$.
@@
By Remark 3.3 and Proposition 2.2, it suffices to consider when $N\leq n-1$ and #$\#J\leq m-1$. Suppose on the contrary that $f$ is oblique. Pick $g,C$ as in Lemma 4.1 with $g$ oblique and denote by $\rho$ the contraction ratio of $g$. If $N=1$, then it is not hard to see that every horizontal slice of $K$ is a finite set. From the form of $C$ (recall (4.5)), $C$ is a finite set. But by Proposition 2.1 and the obliqueness of $g$, $\pi_{1}(g(K))$ is an infinite set. This contradicts $\pi_{1}(g(K))\subset\pi_{1}(C\times\mathbb{R})=C$.
Now suppose $N\geq 2$. By Remark 4.2 (see (4.6)), we can find $z\in\pi_{2}(K)$, $c\in\mathbb{R}$, $k\geq 1$ and a level-$k$ rectangle $R$ such that
@@
Next, since $h_{z}(K^{z})\subset K^{z}$, by Corollary 2.7, $h_{z}$ maps a left isolated point $a$ and a right one $b$ both to one-sided isolated points of $K^{z}$. Since $h_{z}(K^{z})\subset K^{z}$ is simply a one-dimensional reformulation of the inclusion $g\varphi_{R}(K^{z}\times\{z\})\subset L_{z}\cap((K^{z}+c)\times\mathbb{R})$, this in turn gives us two points $\widetilde{a},\widetilde{b}\in g\varphi_{R}(K^{z}\times\{z\})$ and some $\delta>0$ such that both of the tubes $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$ and $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$ do not intersect $(K^{z}+c)\times\mathbb{R}$. Moreover, there are two level-$t$ rectangles $R_{t},R^{\prime}_{t}\subset R$ such that $\widetilde{a}\in g\varphi_{R_{t}}(K)$ and $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, respectively, and (again by Corollary 2.7) $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all $t\geq k_{0}$ for some $k_{0}$. See Figure 5 for an illustration.
@@
Figure 5. An illustration of $\widetilde{a},\widetilde{b},g(R_{t})$ and $g(R^{\prime}_{t})$, where $(K^{z}+c)\times\mathbb{R}$ is supported in the shaded region. (color online)
Pick $x,x^{\prime}\in K$ with $\pi_{2}(x)=\max\pi_{2}(K)$ and $\pi_{2}(x^{\prime})=\min\pi_{2}(K)$. For convenience, let $e_{1}:=(1,0)$ and $e_{2}:=(0,1)$. Since $\widetilde{a}\in g\varphi_{R_{t}}(K)$, $\widetilde{b}\in g\varphi_{R^{\prime}_{t}}(K)$, $x$ can be expressed as
@@
So at least one of $d_{t},-d^{\prime}_{t}$ is no less than $\beta/2$. Writing $e_{1,1}:=\pi_{1}(g(e_{1}))$ and $e_{2,1}:=\pi_{1}(g(e_{2}))$, we have by the obliqueness of $g$ that $e_{1,1},e_{2,1}\neq 0$. We distinguish four simple cases.
Case 1: $d_{t}\geq\beta/2$ and $e_{2,1}>0$
@@
$g\varphi_{R^{\prime}_{t}}(x)$ $=\widetilde{b}+g\varphi_{R^{\prime}_{t}}(\gamma_{t}e_{1})+g\varphi_{R^{\prime}_{t}}(d_{t}e_{2})$
$=\widetilde{b}+n^{-t}\gamma_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|\gamma_{t}|\leq 1$, $0<n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<\delta$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x))=\pi_{1}(\widetilde{b})+n^{-t}\gamma_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta).$
In other words, $g\varphi_{R^{\prime}_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta)\times\mathbb{R}$. But we have seen that this tube does not intersect $(K^{z}+c)\times\mathbb{R}\supset g\varphi_{R}(K)\ni g\varphi_{R^{\prime}_{t}}(x)$, which leads to a contradiction.
Case 2: $d_{t}\geq\beta/2$ and $e_{2,1}<0$
@@
$g\varphi_{R_{t}}(x)$ $=\widetilde{a}+g\varphi_{R_{t}}(c_{t}e_{1})+g\varphi_{R_{t}}(d_{t}e_{2})$
$=\widetilde{a}+n^{-t}c_{t}g(e_{1})+m^{-t}d_{t}g(e_{2}).$
Since $d_{t}\geq\beta/2$, $n>m$ and $|c_{t}|\leq 1$, we have $-\delta<n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}<0$ for all large $t$ and hence
$\pi_{1}(g\varphi_{R_{t}}(x))=\pi_{1}(\widetilde{a})+n^{-t}c_{t}e_{1,1}+m^{-t}d_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})).$
In other words, $g\varphi_{R_{t}}(x)$ lies in the tube $(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a}))\times\mathbb{R}$. Again, we have seen that this tube does not intersect $(K^{z}+c)\times\mathbb{R}\supset g\varphi_{R}(K)\ni g\varphi_{R_{t}}(x)$, which leads to a contradiction.
Case 3: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}>0$
@@
$\pi_{1}(g\varphi_{R^{\prime}_{t}}(x^{\prime}))=\pi_{1}(\widetilde{b})+n^{-t}\gamma^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{b}),\pi_{1}(\widetilde{b})+\delta),$
which leads to a contradiction.
Case 4: $d^{\prime}_{t}\geq\beta/2$ and $e_{2,1}<0$
@@
$\pi_{1}(g\varphi_{R_{t}}(x^{\prime}))=\pi_{1}(\widetilde{a})+n^{-t}c^{\prime}_{t}e_{1,1}+m^{-t}d^{\prime}_{t}e_{2,1}\in(\pi_{1}(\widetilde{a})-\delta,\pi_{1}(\widetilde{a})),$
which leads to a contradiction. ∎
Combining with the above result and Proposition 2.6, we can quickly prove Theorem 1.5.